OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? x = −9, x = 21 x = −6, x = 18 x = 6, x = −18 No solution

OpenStudy (anonymous):

@Hero

hero (hero):

@BrianaNicole16 begin by subtracting 14 from both sides.

OpenStudy (anonymous):

idk how to

hero (hero):

\(2|x − 6| + 14 \color{green}{- 14} = 38 \color{green}{- 14}\)

hero (hero):

@BrianaNicole16 what is 14 - 14 ?

OpenStudy (anonymous):

0 @Hero

hero (hero):

So now we have \(2|x − 6| = 38 \color{green}{- 14}\) @BrianaNicole16, What is \(38 - 14\)?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

@Hero

OpenStudy (lynfran):

2|x-6|=24... now divide both sides by 2..

OpenStudy (anonymous):

how? @LynFran

OpenStudy (lynfran):

2|x-6|/2=24/2 the 2 will cancel frm the left side of the equation leaving |x-6|=24/2

OpenStudy (lynfran):

wats 24/2=?

OpenStudy (anonymous):

12

OpenStudy (lynfran):

so we have |x-6|=12

OpenStudy (lynfran):

now we need to get rid of the absolute value sign...to do that we split the equation into 2... one negative and one positive...like this.. (x-6)=12 and (x-6)= -12

OpenStudy (anonymous):

ok what next

OpenStudy (lynfran):

we solve each equation for x

OpenStudy (anonymous):

ok

OpenStudy (lynfran):

hint:add 6 to both sides of the equation

OpenStudy (lynfran):

so we have x-6+6=12+6 ... and x-6+6= -12+6 x=? x=?

OpenStudy (anonymous):

for x-6+6=12+6 x is 18

OpenStudy (lynfran):

yes and wats the other x=?

OpenStudy (anonymous):

-6

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

thank you so much